OpenStudy (anonymous):

calculus: example indeterminate form lim (f(x) - f(2))/x-2 if f(x) = x^2 - 3x if x approaches 2.

OpenStudy (anonymous):

the prof answer is 1 but i want to learn where did he get solution ((x-2) (x-1)/(x-2) the (x-1)

OpenStudy (nenadmatematika):

f(2)=4-6=-2 so f(x)-f(2)=x^2-3x+2=(x-2)(x-1)..... when you divide it with (x-2) you get x-1 so lim of (x-1) as x approaches to 2 is 1

OpenStudy (anonymous):

ahhhh ic ic so if x approaches to 3 = 2? now i know

OpenStudy (anonymous):

= (x^2-3x - [(2)^2 - 3(2)])/x-2 =lim x approaches 2 (x^2 - 3x - (-2)/x-2 =lim x approaches 2 (x^2 - 3x +2)/x-2 = (4-3(2)+2)/2-2 = 0/0 =lim x approaches 2 ((x-2)(x-1))/(x-2) =lim x approaches 2 (x-1) = 2-1 = 1 answer solution of my prof ^^

OpenStudy (anonymous):

yes i learn again yes im slow but i learn again yes!